) Calculate the mole fraction of the total ions in an aqueous solution prepared by dissolving 0.400 moles of CaF2 in 850.0 g of water.
Answer #1

moles of CaF_2 = 0.400

moles of Ca ions = 0.400

moles of fluoride ions = 0.400 \times 2 = 0.800

moles of water = mass/molecular weight = \frac {850}{18} = 47.23 moles

mole fraction of all ions  = \frac {moles\ of\ CaF_2}{Total\ number\ of\ moles } = \frac {0.4 + 0.8}{0.4 + 47.23} = 0.0248